A Ladder of 6.0 meters in length weighs 150 N rests on horizontal ground and leans at an angle of 65 degrees with the horizontal against a smooth vertical wall. How far up the ladder may a 700 N person go before the ladder slips? The coefficient of friction between ladder and ground is 0.40.
1
Expert's answer
2020-11-10T10:00:19-0500

The diagram:



Assume that the system is in equilibrium when the person is at the highest point "l" meters from the top of the ladder (measured along the ladder).

The force of friction:


"f=\mu N_F."


The equilibrium of forces:


"x:N_W-f=0,\\
y: N_F-W-F=0."

Substitute the force of friction and find the normal forces:


"N_F=W+F,\\
N_W=\mu N_F=\mu(W+F)."

Now, it's time to write the torque around the lower end:


"W\text{ cos}\theta\cdot\frac L2+F\text{ cos}\theta\cdot(L-l)-N_WL\text{ sin }\theta=0,"

substitute the normal force from the wall:


"W\text{ cos}\theta\cdot\frac L2+F\text{ cos}\theta\cdot(L-l)-\\-\mu(W+F)L\text{ sin }\theta=0,\\
l=0.4\text{ m}."

The person may climb 5.6 m before the ladder starts slipping.